OpenStudy (fibonaccichick666):

Inverse Laplace transform question: \[Y(t)=\frac{\frac{1}{s+1}+1}{s^2-2s+2}\] It came from the IVP \(y''-2y'+2y=e^{-t}\) I just cannot seem to reduce it into a doable inverse

OpenStudy (perl):

you have two equal signs in your diff. equation

OpenStudy (fibonaccichick666):

@ganeshie8, @Zarkon, @Mashy Do y'all have any ideas? I've tried partial fractions but it didn't work

OpenStudy (fibonaccichick666):

oops, first should be a minus

OpenStudy (perl):

you can edit your question

OpenStudy (fibonaccichick666):

better?

OpenStudy (fibonaccichick666):

@perl , any ideas on the simplification though?

OpenStudy (anonymous):

\[\begin{align*}Y(s)&=\frac{\dfrac{1}{s+1}+1}{s^2-2s+2}\\\\ &=\frac{\dfrac{1+s+1}{s+1}}{(s-1)^2+1}\\\\ &=\frac{s+2}{(s+1)((s-1)^2+1)}\end{align*}\] Partial fractions from here?

OpenStudy (fibonaccichick666):

yea, that's what I tried, but I got a=0 contradicting with a= -2

OpenStudy (fibonaccichick666):

did I just mess up the algebra?

OpenStudy (anonymous):

\[\begin{align*} \frac{s+2}{\cdots}&=\frac{A}{s+1}+\frac{Bs+C}{(s-1)^2+1}\\\\ s+2&=A(s^2-2s+2)+(Bs+C)(s+1)\\\\ &=(A+B)s^2+(-2A+B+C)s+2A+C \end{align*}\] which gives the system \[\begin{cases} A+B=0\\ -2A+B+C=1\\ 2A+C=2 \end{cases}~~\implies~~A=\frac{1}{5},~B=-\frac{1}{5},~C=\frac{8}{5}\]

OpenStudy (fibonaccichick666):

oh... yep

OpenStudy (fibonaccichick666):

I forgot you have to do a Bs+C...ughhh

OpenStudy (fibonaccichick666):

nooooo wonder

OpenStudy (anonymous):

Mmm yeah it's a quadratic factor in disguise :P

OpenStudy (fibonaccichick666):

I'm good from here. thanks

OpenStudy (anonymous):

You're welcome